The width of a widescreen TV is 10 in. less than its length. The diagonal of the rectangular screen is 10 in. more than the length. Find the length and width of the screen.
How do I set this up and solve?

Answer by mananth(16946)

length =x
width = (x-10 )
diagonal = (x+10)
They form a right triangle.
Apply Pythagoras theorem
(x+10)^2=x^2+ (x-10)^2
x^+20x+100=x^2+x^2-20x+100
2x^2=40x
/x
2x=40
x= 20 the length of the TV